# Optimal upper bounds for anti-canonical volumes of singular toric Fano varieties

### v1

@@ 4. Bounding the dual normalized volumes of Fano simplices
Fix two integers $d\geq 2$ and $q\geq 1$. In this section, we give the upper bound for the dual normalized volume of any $d$-dimensional $\frac{1}{q}$-lc Fano simplex, which corresponds to the upper bound for the anti-canonical volume of a $d$-dimensional $\frac{1}{q}$-lc toric Fano variety with Picard number one. We construct examples to show that the upper bounds are optimal in each dimension $d\geq 2$.
@@
Fix two integers $d\geq 2$ and $q\geq 1$. Then for any $d$-dimensional $\frac{1}{q}$-lc Fano simplex $S$, we have
@@
    if $d\geq 3$, then $\operatorname{Vol}(S^{*})\leq\frac{2u_{d,q}^{2}}{q^{d+1}}$.
@@
So the conclusions follow from Theorem 3.9. ∎
@@
    if $d\geq 3$, then $(-K_{X})^{d}\leq\frac{2u_{d,q}^{2}}{q^{d+1}}$.
@@
By assumption, the associated fan of $X$ is spanned by the faces of a $d$-dimensional $\frac{1}{q}$-lc Fano simplex $S$. Then by [8, Theorem 13.4.3], we have
@@
So we can conclude the proof by Theorem 4.1. ∎
@@
Second, we want to show that $\frac{1}{q}=\max\{t|\operatorname{int}(tS)\cap{\mathbf{N}}=\{{\mathbf{0}}\}\}$. Take a real number $q^{\prime}$ such that $0<q^{\prime}\leq q$. Suppose $v\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbf{N}}$ has barycentric coordinates $\lambda_{1},\cdots,\lambda_{d+1}$, i.e., $\Sigma_{i=1}^{d+1}\lambda_{i}=1$, $\lambda_{i}>0,\forall i\in\{1,\cdots,d+1\}$, and $v=\Sigma_{i=1}^{d+1}\lambda_{i}\frac{1}{q^{\prime}}v_{i}\in{\mathbf{N}}$. Then for $i\in\{1,\cdots,d-1\}$, $\lambda_{i}=\frac{q^{\prime}z_{i}+q}{1+u_{i,q}}$ for some $z_{i}\in{\mathbb{Z}}$, and $\lambda_{d+1}-\lambda_{d}=q^{\prime}z$ for some $z\in{\mathbb{Z}}$.
If $q^{\prime}=q$, then $\lambda_{i}>0$ implies $\lambda_{i}\geq\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. By (1.1) and the assumption that $v$ is an interior lattice point of $\frac{1}{q}S$, we must have $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. And then $\lambda_{d}+\lambda_{d+1}=\frac{q}{u_{d,q}}$, which implies $\lambda_{d}=\lambda_{d+1}=\frac{q}{2u_{d,q}}$. So we conclude that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$.
If $0<q^{\prime}<q$, then we can show that $-e_{d-1}\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbf{N}}$ by taking $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-2\}$, $\lambda_{d-1}=\frac{q-q^{\prime}}{1+u_{d-1,q}}$ and $\lambda_{d}=\lambda_{d+1}=\frac{q^{\prime}}{2(1+u_{d-1,q})}+\frac{q}{2u_{d,q}}$.
@@
$\mathbb{P}(w_{1},\cdots,w_{d+1}),\quad\text{where}\quad w_{i}:=|\det(v_{j}:j\in\{1,\cdots,d+1\}\setminus\{i\})|$
@@
Let $e_{1},\cdots,e_{d}$ be the standard basis of $\mathbb{R}^{d}$. Set $e=\Sigma_{i=1}^{d-1}e_{i}$. Let $S:=\operatorname{conv}((1+u_{1,q})e_{1}-qe,\cdots,(1+u_{d-1,q})e_{d-1}-qe,-e_{d}-qe,e_{d}-qe)$. Then $S^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1}^{*},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1}^{*},\pm\frac{u_{d,q}}{q}e_{d}^{*})$, where $e_{1}^{*},\cdots,e_{d}^{*}$ are the dual basis of $e_{1},\cdots,e_{d}$.
@@
Let $\beta_{1},\cdots,\beta_{d+1}$ be the barycentric coordinates of ${\mathbf{0}}$ with respect to $S$. Then
@@
Set $u_{i}:=\frac{1}{q}e_{i}^{*}$ for $11\leq i\leq d-$ 1 and
$u_{d}:=-\frac{u_{d,q}}{q}e_{d}^{*}-\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j}^{*},u_{d+1}:=\frac{u_{d,q}}{q}e_{d}^{*}-\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j}^{*}.$
We claim that $S^{*}=\operatorname{conv}(u_{1},\cdots,u_{d+1})$. According to the standard duality (see [3, Proposition 6.1] and [14, Proposition 3.6]), it suffices to show that $<(1+u_{i,q})e_{i}-qe,u_{j}>=-1$ for $i\in\{1,\cdots,d-1\}$ with $j\neq i$, $<-e_{d}-qe,u_{j}>=-1$ for $j\neq d$, and $<e_{d}-qe,u_{j}>=-1$ for $j\neq d+1$. So the claim follows directly from the construction.
Set $Q_{1}:=S^{*}+\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j}^{*}$. We need to show that
$Q_{1}\simeq Q_{2}:=\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1}^{*},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1}^{*},\pm\frac{u_{d,q}}{q}e_{d}^{*}).$
Define $\phi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ by sending $e_{i}^{*}$ to $\frac{q}{1+u_{i,q}}(\frac{1}{q}e_{i}^{*}+\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j}^{*})$ for $i\in\{1,\cdots,d-1\}$, and mapping $e_{d}^{*}$ to $e_{d}^{*}$. We claim that $\phi$ is an integral transformation. By (1.2), we have
@@
So we can rewrite $\phi(e_{i}^{*})$ as
$\frac{1+u_{i,q}\Pi_{l=i+1}^{d-1}(1+u_{l,q})}{1+u_{i,q}}e_{i}^{*}+\Sigma_{1\leq j\leq d-1,j\neq i}\frac{u_{d,q}}{(1+u_{i,q})(1+u_{j,q})}e_{j}^{*}$
@@
Set $e^{\prime}:=\Sigma_{j=1}^{d-1}e_{j}^{*}$. Define $\psi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ by sending $e_{i}^{*}$ to $(1+u_{i,q})e_{i}^{*}-qe^{\prime}$ and mapping $e_{d}^{*}$ to $e_{d}^{*}$. Certainly, $\psi$ is an integral transformation and for $i\in\{1,\cdots,d-1\}$, we have
$\psi\circ\phi(e_{i}^{*})$ ${}=\frac{q}{1+u_{i,q}}(\frac{1+u_{i,q}}{q}e_{i}^{*}-e^{\prime}+\frac{u_{d,q}}{q}e^{\prime}-\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e^{\prime})$
${}=e_{i}^{*},$
@@
where the last equality follows from (1.1). So $S^{*}$ is unimodularly equivalent to $\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1}^{*},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1}^{*},\pm\frac{u_{d,q}}{q}e_{d}^{*})$.

### v2

@@ 4. Bounding the dual normalized volumes of Fano simplices
Fix two integers $d\geq 2$ and $q\geq 1$. In this section, we give the upper bound for the dual normalized volumes of $d$-dimensional $\frac{1}{q}$-lc Fano simplices, which corresponds to the upper bound for the anti-canonical volumes of $d$-dimensional $\frac{1}{q}$-lc toric Fano varieties with Picard number one. We construct examples to show that the upper bounds are optimal in each dimension $d\geq 3$.
@@
Fix two integers $d\geq 2$ and $q\geq 1$. Then for any $d$-dimensional $\frac{1}{q}$-lc Fano simplex $S\subset\mathbb{R}^{d}$, we have
@@
    if $d\geq 3$, then $\operatorname{Vol}(S^{*})\leq\frac{2u_{d,q}^{2}}{q^{d+1}}$, and
  3. (c)
    for $q\geq 2$, the equality holds if and only if
    $S^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d}),$
    where $e_{i}$ are the standard basis of $\mathbb{R}^{d}$.
@@
So the conclusions $(a)$ and $(b)$ follow from Theorem 3.9. The proof of conclusion $(c)$ will be presented at the end of this section. ∎
@@
    if $d\geq 3$, then $(-K_{X})^{d}\leq\frac{2u_{d,q}^{2}}{q^{d+1}}$, and
  3. (c)
    for $q\geq 2$ the equality holds if and only if
    $X\simeq\mathbb{P}(\frac{2u_{d,q}}{1+u_{1,q}},\cdots,\frac{2u_{d,q}}{1+u_{d-1,q}},1,1).$
@@
By assumption, the associated fan of $X$ is spanned by the faces of a $d$-dimensional $\frac{1}{q}$-lc Fano simplex $S$. Then by [9, Theorem 13.4.3], we have
@@
So the conclusions follow from Theorem 4.1 along with the following Example 4.3 and the proof of Theorem 4.1$(c)$. ∎
@@
Second, we want to show that $\frac{1}{q}=\max\{t|\operatorname{int}(tS)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}\}$. Take a real number $q^{\prime}$ such that $0<q^{\prime}\leq q$. Suppose $v\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbb{Z}}^{d}$ has barycentric coordinates $\lambda_{1},\cdots,\lambda_{d+1}$, i.e., $\Sigma_{i=1}^{d+1}\lambda_{i}=1$, $\lambda_{i}>0,\forall i\in\{1,\cdots,d+1\}$, and $v=\Sigma_{i=1}^{d+1}\lambda_{i}\frac{1}{q^{\prime}}v_{i}\in{\mathbb{Z}}^{d}$. Then for $i\in\{1,\cdots,d-1\}$, $\lambda_{i}=\frac{q^{\prime}z_{i}+q}{1+u_{i,q}}$ for some $z_{i}\in{\mathbb{Z}}$, and $\lambda_{d+1}-\lambda_{d}=q^{\prime}z$ for some $z\in{\mathbb{Z}}$.
If $q^{\prime}=q$, then $\lambda_{i}>0$ implies $\lambda_{i}\geq\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. By (1.1) and the assumption that $v$ is an interior lattice point of $\frac{1}{q}S$, we must have $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. And then $\lambda_{d}+\lambda_{d+1}=\frac{q}{u_{d,q}}$, which implies $\lambda_{d}=\lambda_{d+1}=\frac{q}{2u_{d,q}}$. So we conclude that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$.
If $0<q^{\prime}<q$, then we can show that $-e_{d-1}\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbb{Z}}^{d}$ by taking $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-2\}$, $\lambda_{d-1}=\frac{q-q^{\prime}}{1+u_{d-1,q}}$ and $\lambda_{d}=\lambda_{d+1}=\frac{q^{\prime}}{2(1+u_{d-1,q})}+\frac{q}{2u_{d,q}}$.
@@
$\mathbb{P}(w_{1},\cdots,w_{d+1}),\quad\text{where}\quad w_{i}:=|\det(v_{j}:j\in\{1,\cdots,d+1\}\setminus\{i\})|.$
@@
Let $e_{1},\cdots,e_{d}$ be the standard basis of $\mathbb{R}^{d}$. Set $e=\Sigma_{i=1}^{d-1}e_{i}$. Let $S:=\operatorname{conv}((1+u_{1,q})e_{1}-qe,\cdots,(1+u_{d-1,q})e_{d-1}-qe,-e_{d}-qe,e_{d}-qe)$. Then $S^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d})$.
@@
Let $\beta_{1}\geq\cdots\geq\beta_{d+1}$ be the barycentric coordinates of ${\mathbf{0}}$ with respect to $S$. Then
@@
Set $u_{i}:=\frac{1}{q}e_{i}$ for $11\leq i\leq d-$ 1 and
$u_{d}:=-\frac{u_{d,q}}{q}e_{d}-\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j},u_{d+1}:=\frac{u_{d,q}}{q}e_{d}-\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j}.$
We claim that $S^{*}=\operatorname{conv}(u_{1},\cdots,u_{d+1})$. According to the standard duality (see [3, Proposition 6.1] and [15, Proposition 3.6]), it suffices to show that $<(1+u_{i,q})e_{i}-qe,u_{j}>=-1$ for $i\in\{1,\cdots,d-1\}$ and $j\in\{1,\cdots,d+1\}\setminus\{i\}$, $<-e_{d}-qe,u_{j}>=-1$ for $j\in\{1,\cdots,d+1\}\setminus\{d\}$, and $<e_{d}-qe,u_{j}>=-1$ for $j\in\{1,\cdots,d+1\}\setminus\{d+1\}$. So the claim follows directly from the construction.
Set $Q:=S^{*}+\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j}$. We need to show that
$Q\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d}).$
Define $\phi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ by sending $e_{i}$ to $\frac{q}{1+u_{i,q}}(\frac{1}{q}e_{i}+\frac{1}{q}\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e_{j})$ for $i\in\{1,\cdots,d-1\}$, and mapping $e_{d}$ to $e_{d}$. We claim that $\phi$ is an integral transformation. By (1.2), we have
@@
So we can rewrite $\phi(e_{i})$ as
$\frac{1+u_{i,q}\Pi_{l=i+1}^{d-1}(1+u_{l,q})}{1+u_{i,q}}e_{i}+\Sigma_{1\leq j\leq d-1,j\neq i}\frac{u_{d,q}}{(1+u_{i,q})(1+u_{j,q})}e_{j}$
@@
Define $\psi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ by sending $e_{i}$ to $(1+u_{i,q})e_{i}-qe$ and mapping $e_{d}$ to $e_{d}$. Certainly, $\psi$ is an integral transformation and for $i\in\{1,\cdots,d-1\}$, we have
$\psi\circ\phi(e_{i})$ ${}=\frac{q}{1+u_{i,q}}(\frac{1+u_{i,q}}{q}e_{i}-e+\frac{u_{d,q}}{q}e-\Sigma_{j=1}^{d-1}\frac{u_{d,q}}{1+u_{j,q}}e)$
${}=e_{i},$
@@
where the last equality follows from (1.1). So $Q$ is unimodularly equivalent to $\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d})$, and hence
$S^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d}).$
Now we come to finish the proof of Theorem 4.1.
Lemma 4.5.
Let $e_{1},\cdots,e_{d}$ be the standard basis of $\mathbb{R}^{d}$. Let $S\subset\mathbb{R}^{d}$ be the lattice simplex $\operatorname{conv}(v_{1},\cdots,v_{d+1})$ with $v_{d+1}={\mathbf{0}}$. Suppose $\operatorname{int}(S)\cap q{\mathbb{Z}}^{d}=\{\Sigma_{i=1}^{d}\frac{qv_{i}}{1+u_{i,q}}\}$. Then $S\simeq\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d,q})e_{d},{\mathbf{0}})$.
Proof.
By assumption, we have $\Sigma_{i=1}^{d}\frac{v_{i}}{1+u_{i,q}}\in{\mathbb{Z}}^{d}$. Then for $1\leq i\leq d$, we conclude $\frac{v_{i}}{1+u_{i,q}}\in{\mathbb{Z}}^{d}$ by the observation that $1+u_{i,q}$ are coprime to each other.
Let $\Lambda:=\frac{v_{1}}{1+u_{1,q}}{\mathbb{Z}}+\cdots+\frac{v_{d}}{1+u_{d,q}}{\mathbb{Z}}$ be the sublattice of ${\mathbb{Z}}^{d}$. We claim that $\det(\Lambda)=1$. Since by construction,
$\{\Sigma_{i=1}^{d}\frac{qv_{i}}{1+u_{i,q}}\}{}$ $\subset((0,1]\frac{qv_{1}}{1+u_{1,q}}+\cdots+(0,1]\frac{qv_{d}}{1+u_{d,q}})\cap q\Lambda$
$\subset((0,1]\frac{qv_{1}}{1+u_{1,q}}+\cdots+(0,1]\frac{qv_{d}}{1+u_{d,q}})\cap q{\mathbb{Z}}^{d}$
$\subset\operatorname{int}(S)\cap q{\mathbb{Z}}^{d}=\{\Sigma_{i=1}^{d}\frac{qv_{i}}{1+u_{i,q}}\},$
we get $((0,1]\frac{v_{1}}{1+u_{1,q}}+\cdots+(0,1]\frac{v_{d}}{1+u_{d,q}})\cap{\mathbb{Z}}^{d}=\{\Sigma_{i=1}^{d}\frac{v_{i}}{1+u_{i,q}}\}$. So the claim holds because
$\det\Lambda=\frac{\det\Lambda}{\det{\mathbb{Z}}^{d}}=|((0,1]\frac{v_{1}}{1+u_{1,q}}+\cdots+(0,1]\frac{v_{d}}{1+u_{d,q}})\cap{\mathbb{Z}}^{d}|$
by [4, Chapter VII, Theorem 2.5] and by noting that
${}|((0,1]\frac{v_{1}}{1+u_{1,q}}+\cdots+(0,1]\frac{v_{d}}{1+u_{d,q}})\cap{\mathbb{Z}}^{d}|$
$=$ ${}|([0,1)\frac{v_{1}}{1+u_{1,q}}+\cdots+[0,1)\frac{v_{d}}{1+u_{d,q}})\cap{\mathbb{Z}}^{d}|.$
Hence $\Lambda={\mathbb{Z}}^{d}$. Then the map defined by sending $\frac{v_{i}}{1+u_{i,q}}$ to $e_{i}$ for $1\leq i\leq d$ is a unimodular transformation. Thus $S\simeq\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d,q})e_{d},{\mathbf{0}})$. ∎
Lemma 4.6.
Let $e_{1},\cdots,e_{d}$ be the standard basis of $\mathbb{R}^{d}$. Suppose $S\subset\mathbb{R}^{d}$ is a $d$-dimensional $\frac{1}{q}$-lc Fano simplex such that the barycentric coordinates of ${\mathbf{0}}$ with respect to $S$ are
$(\beta_{1},\cdots,\beta_{d+1})=(\frac{q}{1+u_{1,q}},\cdots,\frac{q}{1+u_{d-1,q}},\frac{q}{2u_{d,q}},\frac{q}{2u_{d,q}}),$
Then $S\simeq\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d-1,q})e_{d-1},\mp(v^{(d-1)},h))$ for some $v^{(d-1)}\in\{0,1,\cdots,h-1\}^{d-1}$ and some positive integer $h\in{\mathbb{Z}}$.
Proof.
Let $v_{1},\cdots,v_{d+1}$ be the vertices of $S$ associated to the barycentric coordinates $\beta_{1},\cdots,\beta_{d+1}$. Then $\Sigma_{i=1}^{d+1}\beta_{i}v_{i}={\mathbf{0}}.$ So we get
$\frac{q}{1+u_{1,q}}\cdot v_{1}+\cdots+\frac{q}{1+u_{d-1,q}}\cdot v_{d-1}+\frac{q}{2u_{d,q}}\cdot(v_{d}+v_{d+1})={\mathbf{0}}.$
Since $u_{d,q}$ is divided by $q(1+u_{i,q})$ for each $11\leq i\leq d-$ 1, we have
$\frac{v_{d}+v_{d+1}}{2}=-\Sigma_{i=1}^{d-1}\frac{u_{d,q}}{1+u_{i,q}}\cdot v_{i}\in q{\mathbb{Z}}^{d}.$
Now set $v:=-\frac{v_{d}+v_{d+1}}{2}$, $S^{\prime}:=S+v$ and $v_{i}^{\prime}:=v_{i}+v$ for $11\leq i\leq d+$ 1. Denote by $H$ the hyperplane section $\operatorname{conv}(v_{1}^{\prime},\cdots,v_{d-1}^{\prime},{\mathbf{0}})$ of $S^{\prime}$. By assumption that $S$ is a $\frac{1}{q}$-lc Fano simplex, we have $\operatorname{int}(S)\cap q{\mathbb{Z}}^{d}=\{{\mathbf{0}}\},$ which implies
$\operatorname{int}(S^{\prime})\cap q{\mathbb{Z}}^{d}=\{v\}=\{\Sigma_{i=1}^{d-1}\frac{q}{1+u_{i,q}}v_{i}^{\prime}\}$
by noting that $v\in q{\mathbb{Z}}^{d}$. And then $\{v\}\subset\operatorname{int}(H)\subset\operatorname{int}(S^{\prime})$ implies
$\operatorname{int}(H)\cap q{\mathbb{Z}}^{d}=\{\Sigma_{i=1}^{d-1}\frac{q}{1+u_{i,q}}v_{i}^{\prime}\}.$
So Lemma 4.5 applies to $H$. We may assume $H=\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d-1,q})e_{d-1},{\mathbf{0}})$ by doing some unimodular transformation. By construction, $v_{d}^{\prime}+v_{d+1}^{\prime}={\mathbf{0}}$. So we can write $-v_{d}^{\prime}=v_{d+1}^{\prime}=(v^{(d-1)},h)$ for some $v^{(d-1)}\in{\mathbb{Z}}^{d-1}$ and $h\in{\mathbb{Z}}$. We may assume $h>0$ by possibly interchanging $v_{d}^{\prime}$ and $v_{d+1}^{\prime}$. Let $v^{\prime(d-1)}\in\{0,1,\cdots,h-1\}^{d-1}$ such that $v^{\prime(d-1)}\equiv v^{(d-1)}(\mod{h})$. Assume $v^{(d-1)}=(x_{1},\cdots,x_{d-1})$ and $v^{\prime(d-1)}=(x_{1}+c_{1}h,\cdots,x_{d-1}+c_{d-1}h)$ for some integers $c_{1},\cdots,c_{d-1}$. Consider the $d\times d$ matrix $B:=$
$\begin{bmatrix}1&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ c_{1}&c_{2}&\cdots&c_{d-1}&1\end{bmatrix}$
We conclude the proof by the observation that the transformation induced by $B$ is a unimodular transformation preserving $\mathbb{R}^{d-1}\times\{0\}$ which maps $S^{\prime}$ to $\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d-1,q})e_{d-1},\mp(v^{\prime(d-1)},h))$.
Proof of Theorem 4.1$(c)$.
We only need to prove that for $q\geq 2$, $\operatorname{Vol}(S^{*})=\frac{2u_{d,q}^{2}}{q^{d+1}}$ implies $S^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d}).$ Suppose the equality holds. By Theorem 3.9, we see that the equality holds if only if
$(\beta_{1},\cdots,\beta_{d+1})=(\frac{q}{1+u_{1,q}},\cdots,\frac{q}{1+u_{d-1,q}},\frac{q}{2u_{d,q}},\frac{q}{2u_{d,q}}),$
where $\beta_{1},\cdots,\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}$ with respect to $S$. So by applying Lemma 4.6, we have $S\simeq\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d-1,q})e_{d-1},\mp(v^{(d-1)},h))$ for some $v^{(d-1)}\in\{0,1,\cdots,h-1\}^{d-1}$ and some positive integer $h\in{\mathbb{Z}}$. On the other hand, deducing from [15, Proposition 3.6] or following from [3, Proposition 6.2], we have
$\operatorname{Vol}(S)\cdot\operatorname{Vol}(S^{*})=\frac{1}{\Pi_{i=1}^{d+1}\beta_{i}}.$
So by calculation, $\operatorname{Vol}(S)=\frac{2u_{d,q}}{q}=\Pi_{i=1}^{d-1}(1+u_{i,q})\cdot 2h$. Then we must have $h=1$ and hence $S\simeq\operatorname{conv}((1+u_{1,q})e_{1},\cdots,(1+u_{d-1,q})e_{d-1},\mp e_{d})$. So the conclusion follows from Proposition 4.4. ∎